8. Stepladder. In the stepladde long. The tie-rod BD of length 1-3.00 ft is half way up. A person weighing W = 195 lbs climbs d = 7.00 ft along the ladder side AC, as shown. Assume the floor is frictionless and neglect the weight of the ladder. Find (a) the tension T in the tie-rod and (b) the forces exerted on the ladder by the floor. ure , both

Please be as detailed as you can be
Answer #1

a/D = 3110 心· ベs 2. 1 o 2. Ν= 346,961-N ση sino =